### v1

# The viscoelastic rheology of transient diffusion creep

Abstract
Polycrystalline materials have a viscoelastic rheology where the strains produced by stresses depend on the timescale of deformation. Energy can be stored elastically within grain interiors and dissipated by a variety of different mechanisms. One such dissipation mechanism is diffusionally-accommodated/-assisted grain boundary sliding, also known as transient diffusion creep. Here we detail a simple reference model of transient diffusion creep, based on finite element calculations with simple grain shapes: a regular hexagon in 2D and a tetrakaidecadedron in 3D. The linear viscoelastic behaviour of the finite element models can be well described by a parameterised extended Burgers model, which behaves as a Maxwell model at low frequencies and as an Andrade model at high frequencies. The parametrisation has a specific relaxation strength, Andrade exponent and Andrade time. The Andrade exponent depends only on the angles at which grains meet at triple junctions, and can be related to the exponents of stress singularities that occur at triple junctions in purely elastic models without diffusion. A comparison with laboratory experiments shows that the simple models here provide a lower bound on the observed attenuation. However, there are also clearly additional dissipative processes occurring in laboratory experiments that are not described by these simple models.
1 Introduction
A fundamental question of material science is to understand how a material deforms when it is stressed, i.e. what is the rheology of a material? There are two main approaches to improving our understanding of rheology: one is through careful laboratory experiments, the other is through building models of physical processes that occur within materials. It is only through modelling work, and through understanding the fundamental physics, that results from experiments can be safely extrapolated into parameter regimes beyond those of the laboratory. This is of particular concern in Earth sciences where the length scales and time scales of the laboratory are so small compared to those relevant for the planet and its long history.
The aim of the present manuscript is to provide a simple grain-scale model of rock viscoelasticity that can be used as a basis for interpreting and extrapolating laboratory experiments. The model describes the effective viscoelastic behaviour which arises from the grain-scale diffusion of matter by physical mechanisms referred to as diffusionally-assisted or diffusionally-accommodated grain-boundary sliding [1, 2]; mechanisms also referred to as transient diffusion creep. There is a long history of models of transient diffusion creep, with early pioneering work by Lifshits, Raj and Ashby [3, 4, 5]. The present work builds most closely on the more recent study by Lee et al. [6], who developed a two-dimensional finite element model of a bicrystal consisting of two elastic grains with a grain boundary between. The grain boundary allows for both sliding and diffusion of matter. Here, we model the same physics as Lee et al. [6] but in a different geometry, considering an infinite tiling of hexagonal grains in two-dimensions and a tessellation of tetrakaidecahedral grains in three-dimensions, in turn building on models for steady-state diffusion creep described in [7, 8]. A homogenisation approach is used to determine the effective upscaled viscoelastic tensors which describe the macroscale rheology arising from the grain-scale physics.
The manuscript is organised as follows. The next section describes the model, and results for hexagonal grains are given in section 3 and results for tetrakaidecahedral grains in section 4. This is followed by comparison to laboratory data and previous studies in section 5 and conclusions in section 6. A series of appendices provide the mathematical basis for the modelling work.
2 Model description
The model geometry consists of a tessellation of identical grains. The interior of grains are assumed to deform linear elastically, with intrinsic shear modulus $\mu$ and Poisson’s ratio $\nu$. Grains are assumed to slide freely at the grain boundaries (zero shear stress). Vacancies and atoms can diffuse along the grain boundaries due to gradients in normal stress, leading to the plating out or removal of material along the grain boundaries. Energy is stored elastically within the grain interiors and dissipated by grain boundary diffusion. The medium thus behaves viscoelastically. The linear viscoelastic response of a material can be described in a number of different ways, but one of the most useful is as a function of frequency $\omega$. The strain response $\epsilon(t)$ of a viscoelastic material to a time harmonic shear stress $\sigma(t)=\sigma_{0}\mathrm{e}^{i\omega t}$ is
$\epsilon(t)=J^{*}(\omega)\sigma(t)$ (1)
where $J^{*}(\omega)$ is the complex shear compliance. The reciprocal $G^{*}(\omega)=1/J^{*}(\omega)$ is known as the complex shear modulus. As a complex quantity, $J^{*}(\omega)$ can be split into real and imaginary parts as $J^{*}(\omega)=J_{1}(\omega)-iJ_{2}(\omega)$, where $J_{1}$ is termed the storage compliance and $J_{2}$ is termed the loss compliance. Finite element analysis allows the calculation of the function $J^{*}(\omega)$ for a given grain geometry, which is the main outcome of the present study. A detailed account of the governing equations and their discretisation by the finite element method can be found in appendices A to F.
3 Hexagonal grains
In general the viscoelastic response of a material is described by a fourth-rank complex compliance tensor, which in three-dimensions involves an array of 81 complex numbers as a function of frequency. However, the number of independent components is greatly reduced on consideration of symmetry, both in the governing equations and in the geometry. In 2D, a tiling of regular hexagons yields fourth rank tensors which are isotropic, i.e. there is no dependence on orientation, and the viscoelastic behaviour can be described by just two numbers as a function of frequency: a complex bulk compliance and a complex shear compliance. The assumed grain scale physics does not allow for any dissipation under bulk deformation and so the only viscoelastic behaviour that need be explicitly computed is that under shear.
Let us consider the end-member behaviour at zero and infinite frequency. At zero frequency (infinite period) the diffusion process reaches a steady-state. This steady-state deformation is steady-state diffusion creep (Coble creep [9]) and the material behaves as a Newtonian viscous fluid. The steady-state shear viscosity $\eta$ for an array of hexagonal grains has been calculated by many authors [e.g. 10, 11, 7] and is
$\eta=\frac{1}{144}\frac{kTd^{3}}{\delta D_{gb}\Omega}$ (2)
where $k$ is Boltzmann’s constant, $T$ is temperature, $d$ is the distance between opposite sides of the hexagon, $\delta$ is the grain boundary width, $D_{gb}$ is the self-diffusion coefficient for grain boundary diffusion, and $\Omega$ is the atomic volume.
At the other extreme, at infinite frequency, diffusion can be neglected and the behaviour is that of an elastic solid. However, the effective shear modulus is less than the intrinsic shear modulus $\mu$ of the grains because we assume shear stresses are zero on the grain boundaries. We denote this effective modulus as $G_{0}$ and refer to it as the unrelaxed modulus, as there has been no relaxation of the stresses normal to the grain boundaries by the process of main interest here, by grain boundary diffusion.
Figure 1: Properties of the model for hexagonal grains as a function of Poisson’s ratio $\nu$. a) Scaled unrelaxed shear modulus $G_{0}/\mu$ and scaled steady-state shear modulus $G_{ss}/\mu$. b) Relaxation strength $\Delta$ associated with grain boundary diffusion, $\Delta=(G_{0}/G_{ss})-1$. c) Ratio of the Andrade time $\tau_{A}$ to the Maxwell time $\tau_{M}$. The data in these plots can be well fit by simple rational functions of $\nu$ (see Appendix J). The plots show the behaviour for the full allowable range of Poisson’s ratio for linear elasticity, $-1\leq\nu\leq\tfrac{1}{2}$; a typical value for rock is $\nu\sim 0.3$.
Figure 1a shows $G_{0}/\mu$ as a function of Poisson’s ratio $\nu$. This quantity was calculated previously for hexagonal grains by Gharemani [12]; the results here are in complete agreement for the range of Poisson’s ratios used in that study (see Figure 4 of [12]). $G_{0}/\mu$ ranges from 0.6 to 0.86 depending on Poisson’s ratio. For a typical value of Poisson’s ratio of rock of around $\nu=0.3$, the ratio $G_{0}/\mu=0.81$.
The ratio of the steady-state viscosity $\eta$ to the unrelaxed modulus $G_{0}$ defines a timescale
$\tau_{M}=\frac{\eta}{G_{0}}$ (3)
which we refer to as the Maxwell time. By scaling the governing equations it can be shown that the Maxwell time is the only natural timescale in the problem, and hence it is useful to present all results in terms of this timescale, i.e. in terms of a dimensionless frequency $\omega\tau_{M}$.
Figure 2: A summary of the viscoelastic response for hexagonal grains with a Poisson’s ratio $\nu=0.3$. a) $|G^{*}|/G_{0}$, the ratio of the absolute shear modulus to the unrelaxed modulus as a function of dimensionless frequency $\omega\tau_{M}$. b) $Q^{-1}$, inverse quality factor (attenuation) as a function of dimensionless frequency. c) Scaled compliances $(G_{0}J_{1}-1)/\Delta$ and $(G_{0}J_{2}-1/\omega\tau_{M})/\Delta$ as a function of dimensionless frequency. d) Cole-Cole plot of the scaled compliances against each other. In all plots solid lines show the finite element solutions, dashed lines an extended Burgers model fit, dotted lines an Andrade model fit, dash-dotted lines a Maxwell model fit
Figure 2 shows plots which summarise the viscoelastic behaviour of the model for hexagonal grains. The top two panels show quantities commonly plotted in experimental studies of forced oscillation; the bottom two panels show scaled versions of the compliances which illustrate the nature of the relaxation. Figure 2 also shows fits to the finite element results using some simple viscoelastic models: the Maxwell model, the Andrade model and the extended Burgers Model (see Appendix H for a description of these models).
The behaviour of the model is best understood by considering its extremes. As remarked earlier, the behaviour at zero frequency is that of a viscous fluid. The behaviour at low frequencies $\omega\tau_{M}\ll 1$ more generally (the diffusionally-accommodated regime) is that of a Maxwell model with a dashpot of viscosity $\eta$ in series with a spring of compliance $J_{ss}$ or modulus $G_{ss}=1/J_{ss}$. $\eta$ is the steady-state viscosity and $G_{ss}$ is the steady-state modulus. The steady-state compliance $J_{ss}$ is the limiting value of the storage compliance: $J_{1}\rightarrow J_{ss}$ as $\omega\tau_{M}\rightarrow 0$. The relaxation strength $\Delta$ is the fractional increase in compliance due to the relaxation of stresses by grain boundary diffusion, given by
$\Delta=\frac{J_{ss}}{J_{0}}-1=\frac{G_{0}}{G_{ss}}-1.$ (4)
The relaxation strength is plotted in Figure 1b and varies with Poisson’s ratio $\nu$ from $\Delta=0.70$ to 0.25, with a value of $\Delta=0.33$ for a typical rock value of $\nu=0.3$.
At high frequencies (the diffusionally-assisted regime), $\omega\tau_{M}\gg 1$, the behaviour is that of an Andrade model, with power-law variation of attenuation as $Q^{-1}\propto\left(\omega\tau_{M}\right)^{-\alpha}$ for a particular exponent $\alpha$. The Andrade exponent $\alpha$ can be calculated explicitly using singularity analysis [13, 14, 15] following a scaling argument due to Lee et al. [6]. In the absence of grain boundary diffusion, the pure elastic problem with freely-sliding grain-boundaries has singularities in the stress at the triple junctions where three grains meet, i.e. the normal stresses go as $\sigma_{nn}\propto r^{-\lambda}$ where $r$ is the distance from the triple junction. The analysis of Picu and Gupta [13] shows that for a triple junction with equal angles of 120° and freely sliding grain boundaries the exponent $\lambda=0.4491862$.
In the presence of diffusion, the singularity in the stress is removed. The length scale over which diffusion acts depends on the frequency of the forcing and scales with
$l=\left(\frac{\mu\delta D_{gb}\Omega}{kT\omega}\right)^{1/3}.$ (5)
Provided $l\ll d$ the dissipation due to grain boundary diffusion is concentrated in a small neighbourhood of the triple junction, whereas elastic energy is stored throughout the grain interior. Scaling analysis shows that the power law exponent $\alpha$ in the Andrade model is related to the stress singularity exponent $\lambda$ as
$\alpha=\frac{2}{3}\left(1-\lambda\right).$ (6)
For a derivation see Appendix K and section 4 of [6]. From the Picu and Gupta result [13], the exponent $\alpha=0.3672092$.
The power law behaviour can be clearly seen in Figure 2, both in the high frequency end of the attenuation plot in panel b, but also in the scaled compliance plots in panels c and d. In panel c the power-law behaviour manifests as heavy tails at high frequencies, which can be contrasted with a simple Debye peak one would get for a Burgers rheology with a single relaxation time. Similarly, the behaviour near the origin in the Cole-Cole plot of panel d is tangent to a line with slope $\tan\pi\alpha/2$, quite different from a Burgers rheology which would plot as a semicircle on a Cole-Cole diagram.
The other parameter which specifies the Andrade model is the Andrade time $\tau_{A}$ (appendix H.4). Unlike the exponent $\alpha$, this time cannot be determined purely by examining the singular behaviour at the triple junctions, but is also sensitive to the overall grain geometry and Poisson’s ratio. Figure 1c plots the Andrade time calculated from the finite element analysis as a function of Poisson’s ratio, and shows that the Andrade time is broadly similar to the Maxwell time, increasing from $\tau_{A}/\tau_{M}=0.36$ to 2.13. For a typical rock value of $\nu=0.3$, $\tau_{A}/\tau_{M}=1.315$.
The Maxwell model and the Andrade model characterise the behaviour of the finite element results at low and high dimensionless frequency respectively. Also shown in Figure 2 is another model fit to the finite element data using an extended Burgers Model (appendix H.5). This model is similar to the Andrade model in having a relaxation spectrum with a power-law distribution of relaxation times, but has a long-period cut-off to the relaxation spectrum given by a parameter $\tau_{H}$ which can be related to $\tau_{A}$ and $\Delta$ by $\tau_{H}=\tau_{A}\left(\Gamma(1-\alpha)\Delta\right)^{1/\alpha}$, where $\Gamma(z)$ is the Gamma function. The extended Burgers model has the same asymptotic behaviour as the finite element model at both low and high frequencies. As can be seen in Figure 2, the extended Burgers model provides an excellent fit to the finite element results, with only subtle differences around $\omega\tau_{M}=20$ to be seen in panels a and b which show absolute shear modulus and attenuation. The differences are somewhat more noticeable in the plots of scaled compliance in panels c and d, but even in these plots the differences are slight. Thus an effective parametrisation of the finite element results is to use an extended Burgers model with the appropriate values of $G_{0}$, $\tau_{M}$, $\Delta$ (Figure 1b), $\alpha=0.3672092$ and $\tau_{A}/\tau_{M}$ (Figure 1c).
4 Tetrakaidecahedral grains
Figure 3: Images of boundary normal stress $\sigma_{nn}$ for time-harmonic deformation of frequency $\omega$ for tetrakaidecahedral grains. The imposed macroscale strain is proportional to $\cos\omega t$ where $t$ is time, and the principal axes of the deformation are aligned with the square faces which are at right angles to each other. Orange colours are positive, blue colours are negative, and off-white is zero. Left column shows images at low dimensionless frequency $\omega\tau_{M}=0.1$; right column shows images at high dimensionless frequency $\omega\tau_{M}=500$. The top row is at time $t=0$ in the cycle, the bottom row at a quarter of a period later. The bottom left image is similar to that shown for steady-state creep in Figure 9 (networked) in [16]. At high dimensionless frequency diffusion only influences a narrow boundary layer in the neighbourhood of the triple line (the lengthscale $l$ in (5)); at low dimensionless frequency diffusion influences the whole grain boundary.
A simple grain shape suitable for 3D calculations is the tetrakaidecadedron (truncated octahedron). This grain shape can tessellate three-dimensional space, and has 14 faces, 6 of which are squares and 8 of which are hexagons (Figure 3). Unlike the hexagonal geometry in 2D, for which any fourth rank tensor (like the compliance tensor) must be isotropic, the fourth rank tensor associated with a tessellation of tetrakaidecahedrons is in general anisotropic, although there are strong constraints provided by the cubic symmetry. With cubic symmetry the resistance to shear is described by two independent parameters.
The steady-state Coble creep viscosity tensor of a tiling of tetrakaidecahedrons was calculated in [7] and expressed in terms of two shear viscosities $\eta^{(1)}$ and $\eta^{(2)}$,
$\eta^{(1)}=0.0012242\frac{kTd^{3}}{\delta D_{gb}\Omega},\quad\eta^{(2)}=0.0034715\frac{kTd^{3}}{\delta D_{gb}\Omega},$ (7)
where $\eta^{(1)}$ is the shear viscosity associated with deformation whose principal axes align with the square faces of the grain, and $\eta^{(2)}$ is associated with shear deformation at an angle of 45° to those axes. $d$ is defined as the distance between opposite square faces. One way of defining an average viscosity is in terms of a Voigt-average,
$\eta^{V}=\frac{2}{5}\eta^{(1)}+\frac{3}{5}\eta^{(2)}=0.0025726\frac{kTd^{3}}{\delta D_{gb}\Omega}.$ (8)
Figure 4: The same properties as Figure 1, but for a tessellation of tetrakaidecahedrons in 3D. The properties of such a tiling are anisotropic. Shaded bands are used to show how the parameters can vary with different orientations. The solid lines show results from Voigt-averaging the corresponding complex moduli.
Figure 5: The same viscoelastic response functions as Figure 2 but for the case of tetrakaidecahedral grains. The Poisson’s ratio $\nu=0.3$ and the curves are shown for Voigt-averaged values.
The general viscoelastic response in shear can similarly be expressed in terms of two complex shear moduli as a function of frequency, or in terms of a particular average. Figure 4 presents the same model properties as given in Figure 1, but now for the case of tetrakaidecahedral grains. Due to the anisotropy, the results are presented in terms of bands of values, where the extremes are calculated in terms of the complex moduli associated with the directions for $\eta^{(1)}$ and $\eta^{(2)}$ above. The solid lines show the corresponding Voigt-average values (Appendix I). Figure 5 shows a similar plot to Figure 2 for the case of tetrakaidecahedrons.
The overall viscoelastic response for the tetrakaidecahedral grains is very similar to that for hexagonal grains, with small differences that can be expressed in terms of the different effective values of $\Delta$, $\tau_{A}/\tau_{M}$ and $\alpha$. The differences in $\alpha$ can be attributed to the differences in the angles at which grain boundaries meet at the triple lines. For a hexagon, the three angles are all 120°; for a tetrakaidecadedron the angles are 125.26°-125.26°-109.47° ($125.26\text{\textdegree}=\operatorname{arccos}(-1/\sqrt{3})$, $109.47\text{\textdegree}=\operatorname{arccos}(-1/3)$). The Picu and Gupta [13] stress exponent for these angles is $\lambda=\tfrac{1}{2}$ and hence $\alpha=\tfrac{1}{3}$, a little less than the value for hexagons. For a typical rock value of $\nu=0.3$, the relaxation strength $\Delta=0.55$ and the scaled Andrade time is $\tau_{A}/\tau_{M}=1.7$. The most notable difference between Figure 2 and Figure 5 is due to the change in relaxation strength, which is larger by a factor of 1.7 in Figure 5. The high frequency behaviour is very similar as both the Andrade exponent and the ratio of Andrade time to Maxwell time are similar for the two geometries.
5 Discussion
$\alpha$ $\tau_{A}/\tau_{M}$ $\Delta$ Reference
0.3672092 1.31 0.330 This study, hexagon model, $\nu=0.3$
0.3333333 1.68 0.552 This study, tetrakaidecahedron model, $\nu=0.3$
0.3 11.1 0.202 Lee et al. [6] bicrystal model, sawtooth (S), $\varphi=30\text{\textdegree}$
0.3672092 248 0.0272 Lee et al. [6], bicrystal model, truncated sawtooth (TS), $\varphi=60\text{\textdegree}$
0.38 0.415 - Takei et al. [17], borneol
0.38 0.144 - Yamauchi and Takei [18], borneol + diphenylamine 40, 41, 43
0.33 0.58 1.4 Jackson and Faul [19], synthetic olivine 6585
0.274 0.124 1.04 Jackson and Faul [19], synthetic olivine 6381, 6585, 6365, 6261, 6328
0.26 0.218 1.2 Jackson [20], synthetic olivine
0.217 1.63 0.76 Qu et al. [21], synthetic dunite (SS-jacketed) A1802
0.206 0.0299 1.76 Barnhoorn et al. [22], synthetic MgO 1096 and 1077
0.247 0.216 - Priestley et al. [23], fit to multiple data sets, $T_{h}=0.89$
0.228 0.0269 - Priestley et al. [23], fit to multiple data sets, $T_{h}=1.0$
Table 1: A comparison of the viscoelastic parameters determined in this study with some of those in the literature. $\alpha$ is the Andrade exponent, $\tau_{A}/\tau_{M}$ is the ratio of the Andrade time to the Maxwell time, and $\Delta$ is the relaxation strength. The synthetic olivine/ synthetic dunite samples also contain a small fraction of pyroxene grains. The fits of Priestley et al. [23] are functions of homologous temperature $T_{h}$, the ratio of the absolute temperature to the solidus temperature.
Figure 6: Comparison of the model presented here with literature estimates of a) scaled modulus and b) attenuation. Black lines show the model for tetrakaidecahedrons (solid line) and hexagons (dotted line) for a Poisson’s ratio $\nu=0.3$. Results for the Lee et al. [6] bicrystal model are shown as grey lines: the sawtooth (S) interface as a solid line (from their Figure 6), the truncated sawtooth (TS) interface as a dashed line (from their Figure 7). Yellow dots show borneol samples 27, 32, and 36 from Takei [17]; red dots show borneol + diphenylamine samples 41, 42, and 50 from Yamauchi and Takei [24]. Shading is used to indicate temperature, with darker colours indicating high temperature (see Figure C1 and text of [24] for further details). Green dots show data for synthetic olivine polycrystal sample 6585 of Jackson and Faul [19] for temperatures between 1000 °C and 1200 °C. Purple dots show data for synthetic dunite polycrystal sample A1802 of Qu et al. [21] between 1000 °C and 1200 °C. Estimates of Maxwell time and unrelaxed modulus used for scaling laboratory data use those values given in the studies concerned.
Table 1 shows a comparison between the viscoelastic parameters calculated for the models here with those given in the literature, which have principally been obtained by fits to laboratory experiments. Also given in Table 1 are parameters inferred from the bicrystal model of Lee et al. [6], determined by fitting the curves shown in their figures 6 and 7. Figure 6 provides a comparison between the models and laboratory data in plots of attenuation and scaled modulus.
The bicrystal model of [6] describes the same physics as considered here, and differences only arise due to differences in geometry. Lee et al. [6] consider two types of interface between the two grains of the bicrystal: i) a sawtooth interface (S) and ii) a truncated sawtooth (TS) interface. These interfaces mimic those seen in an array of regular hexagons (referred to as the mode 2 and mode 1 directions in [4]). One important difference between the bicrystal model and the present study is that in the bicrystal model each corner involves two sliding surfaces coming together, whereas at the triple junctions in the present study three sliding surfaces come together. The stress exponents are thus different in general (see [13] Figure 5). For a 120° interior angle with two sliding surfaces there are two values for $\alpha$; a value of 0.3672092 for eigenfields which are symmetric (which is relevant to the TS model, and identical to the array of hexagons here) and a value of 0.3 for eigenfields which are antisymmetric (relevant to the S model).
A more significant difference between the present study and that of Lee et al. [6] is the values of $\tau_{A}/\tau_{M}$ which are significantly larger for the bicrystal models of [6] than the models here; that is to say the bicrystal models of [6] show significantly less attenuation for the same dimensionless frequency (the grey lines are lower than the black lines on Figure 6b). This difference must arise from the overall differences in grain geometry, where it should be noted that the bicrystal model has two lengthscales which define it rather than just one for model here: there is one lengthscale associated with the variations of the interface, and another associated with the width of the bicrystal. Lee et al. [6] present results where the width is 5 times that of the scale of variations on the interface; it is expected that the ratio of these two lengthscales plays an important role in determining the amount of attenuation, as dissipation only takes place on the boundary but elastic energy is stored throughout the grains.
A number of laboratory studies have aimed to characterise the viscoelastic behaviour of rocks by fitting Andrade and extended Burgers models to forced oscillation data. The results of the present study are remarkably similar to those of Takei [17] for polycrystalline borneol, which is an organic compound used as a rock analogue, for which laboratory experiments can be performed close to room temperature. Both the exponent $\alpha$ and the ratio $\tau_{A}/\tau_{M}$ are fairly close between the borneol studies and the model. Takei [17] (and the subsequent Yamauchi and Takei [18, 24] studies) use an Andrade model for fitting the data, and so the total relaxation strength has not been estimated. As is the case in many experimental studies, data is not available at low enough frequencies (i.e. $\omega\tau_{M}\ll 1$) to accurately determine the relaxed modulus. In Figure 6b which shows attenuation, the yellow dots corresponding to the Takei [17] study lie just above the model curve for tetrakaidecahedrons, demonstrating their similarity, particularly at low temperatures.
There are more significant differences between the present model and those based on laboratory experiments using synthetic rock aggregates deformed at high temperature [19, 20, 22, 21]. These studies provide evidence for somewhat lower values of the Andrade exponent $\alpha$ (as low as 0.2). The studies also provide estimates of total relaxation strength $\Delta$ as they are fit with an extended Burgers model, although such estimates are not well-constrained due to the lack of low frequency data. The high frequency data are expected to constrain the Andrade parameters ($\alpha$ and $\tau_{A}$) and not $\Delta$. In the studies of [19, 22, 20, 21] the parameters of the extended Burgers model are not independently chosen: In [19] the high period cutoff is fixed at a specific value of $10^{6}$ s, and in [22, 20, 21] an assumption is made that the high period cutoff is at the Maxwell time, i.e. $\tau_{H}=\tau_{M}$. This assumption constrains the parameters in Table 1 for these studies to satisfy $\tau_{A}/\tau_{M}=\left(\Gamma(1-\alpha)\Delta\right)^{-1/\alpha}$. Note that in the extended Burgers models fits for the hexagons $\tau_{H}=0.166\tau_{M}$ and for the tetrakaidecahedron $\tau_{H}=0.685\tau_{M}$, so the assumption that $\tau_{H}=\tau_{M}$ may be questioned.
Figure 7: Andrade exponent $\alpha$ as a function of junction angle $\theta$ for a triple junction with angles $\theta$, $2180\text{\textdegree}-\theta/$ 2 and $2180\text{\textdegree}-\theta/$ 2. Dotted lines indicate the junction angles for the tetrakaidecahedron ($\theta=109.47\text{\textdegree}$, $\alpha=\tfrac{1}{3}$) and the hexagon ($\theta=120\text{\textdegree}$, $\alpha=0.3672092$).
It is not clear why some laboratory studies show lower Andrade exponents than the models here. In the models, the Andrade exponent is purely controlled by the triple junction geometries, and significant departures from the symmetric 120°-120°-120° junction angles are needed to produce Andrade exponents of 0.2. This is illustrated in Figure 7, which plots the Andrade exponent expected for asymmetric junctions with one angle of $\theta$ and two identical angles of $2180\text{\textdegree}-\theta/$ 2. An exponent of 0.2 could be produced by having a $60\text{\textdegree}-150\text{\textdegree}-150\text{\textdegree}$ junction, but this seems far from what is observed in the microstructures. There may be some other source of anisotropy or geometrical feature that influences the exponent in some experiments, or there may be another dissipation mechanism producing attenuation that is mapped by the fitting to produce a lower effective Andrade exponent.
One additional mechanism of attenuation is that associated with elastically-accommodated grain boundary sliding. No explicit calculations of this mechanism have been made in the present work, but this has been considered in previous studies, notably for hexagons by Ghahremani [12] and in the bicrystal model by Lee et al. [25, 6]. Instead of assuming the grain boundaries are inviscid, one assumes there is a certain grain boundary viscosity $\eta_{gb}$ that describes the resistance to sliding. This grain boundary viscosity leads to an attenuation peak at a particular period $\tau_{e}$, which up to a dimensionless prefactor is at
$\tau_{e}=\frac{\eta_{gb}d}{\mu\delta}.$ (9)
The relaxation strength for this mechanism can be read off from Figure 1a and Figure 4a as
$\Delta_{e}=\frac{\mu}{G_{0}}-1$ (10)
For $\nu=0.3$ the hexagon model yields $\Delta_{e}=0.233$ and the tetrakaidecadedron model yields $\Delta_{e}=0.310$. Ghahremani [12] have shown that the viscoelastic response of this mechanism is well-described by a Burgers model with relaxation at a single period. Somewhat smaller relaxation strengths were found for the bicrystal model of Lee et al. [6, 25] with $\Delta_{e}=0.098$ for the S model and $\Delta_{e}=0.080$ for the TS model.
The relevance and importance of elastically-accommodated grain-boundary sliding in experimental studies is debated. Many studies fit observations by including a high frequency attenuation peak, which some ascribe to elastically-accommodated grain boundary sliding (e.g. [21]) and others do not (e.g. [17, 24]). As noted by Takei [17] (their equation (22)) and Jackson [26] (their equation (13)), a simple model of grain boundary viscosity by Ashby [27] would suggest that the ratio $\tau_{e}/\tau_{M}$ should be around $b^{2}/d^{2}$ where $b\sim 10^{-10}$ m is the atomic size and $d$ is the grain size. For 1 $\mu$m grains, $\tau_{e}/\tau_{M}\sim 10^{-8}$; for 10 $\mu$m grains, $\tau_{e}/\tau_{M}\sim 10^{-10}$; and for 1 mm grains $\tau_{e}/\tau_{M}\sim 10^{-14}$. Thus any peak from elastically-accommodated grain boundary sliding needs data at very high frequencies relative to the Maxwell time to be resolved. The peak that Qu et al. [21] report has $\tau_{e}/\tau_{M}=7\times 10^{-6}$ and a relaxation strength $\Delta_{e}=0.012$. This is a significantly lower relaxation strength than the simple models here predict. The differences in relaxation strength might be attributed due to geometrical factors not modelled here, such as heterogeneity in grain size in real materials [25, 28].
The studies of [17, 18, 24] on borneol have suggested there is a significant high frequency dissipation peak around $\tau_{P}/\tau_{M}=6\times 10^{-5}$ that cannot be attributed to elastically-accommodated grain boundary sliding because of the independence of the peak position on grain size [17]. The peak has been modelled as scaling on the Maxwell time in the same way as for diffusionally-assisted/accommodated grain boundary sliding, but with a peak width and height that varies with homologous temperature. In particular the relaxation strength $\Delta_{p}$ varies from $\Delta_{p}=0.1$ at a homologous temperature of $T_{h}=0.92$ to $\Delta_{p}=0.5$ at a homologous temperature of $T_{h}=1.0$. The evidence for this peak is seen in the attenuation data in Figure 6 as an increase in attenuation with increasing homologous temperature (the darker red dots). There is as yet no micromechanical model which explains this peak, which has tentatively been suggested to arise from changes associated with premelting of grain boundaries as the material approaches the melting point [18, 24]. A notable challenge for future micromechanical models is to explain and quantify this peak.
Understanding the viscoelastic behaviour of rocks at high homologous temperatures is of particular relevance to Earth sciences when interpreting the significant variations in seismic wavespeed, as well as increased attenuation, that occur near the lithosphere-asthenosphere boundary [e.g. 29, 23]. The models here do not produce enough attenuation to explain what is seen in seismology. For example, with a steady-state asthenospheric shear viscosity $\eta=2\times 10^{19}~{}$a s and an intrinsic shear modulus $\mu=70$ GPa, the Maxwell time $\tau_{M}=3\times 10^{8}$ s, i.e. around 10 years. For a surface wave with a period around $T=50~{}$s, the dimensionless frequency is $\omega\tau_{M}=2\pi\tau_{M}/T=4\times 10^{7}$, for which the models yield $Q^{-1}\sim 10^{-3}$ (Figure 6), significantly less than the asthenospheric $Q^{-1}\sim 10^{-2}$ estimated by Ma et al. [29] and given in PREM [30]. Previous work in extrapolating laboratory experiments by Jackson and Faul [19] suggested that diffusionally-accommodated grain boundary sliding can account for the $Q^{-1}\sim 10^{-2}$ level of attenuation at seismic frequencies. This extrapolation was based on a proposed variation of Maxwell time as linear with grain size, $\tau_{M}\propto d$, rather than $\tau_{M}\propto d^{3}$ as expected from theory for grain-boundary diffusion. More recent experiments in the same laboratory [31] have since observed a much stronger grain size dependence with $\tau_{M}\propto d^{3.17}$, closer to the behaviour predicted by theory, and for which extrapolation does not yield sufficient attenuation at seismic frequencies. Thus additional mechanisms of dissipation, such as those associated with premelting [18, 24], impurities [32], elastically-accommodated grain boundary sliding [33], or dislocations [34] are needed to explain the seismic observations.
There a number of natural ways the present modelling work can be extended. One natural extension is to consider diffusion within the interior of grains rather than along the grain boundaries. An example of such a calculation for hexagons is given in Appendix L, and the most notable result is that the relationship between the Andrade exponent and the stress singularity exponent changes to $\alpha=1-\lambda$, such that $\alpha=1/2$ is expected for tetrakaidecahedral grains. Calculations could be performed using a synthetic rock texture which has a range of grain sizes, e.g. using Neper [35]. The expectation would be that the Andrade exponent would remain determined by the triple junction angles, but that the effective $\tau_{A}/\tau_{M}$ and $\Delta$ parameters may change. All the calculations here are for materials in solid-state; solid both in the grain interiors and on the grain boundaries. The effect of melt and melt geometry [16] could be explored, as was done for the case of steady-state diffusion creep viscosity in [7]. The presence of melt allows for a steady-state bulk viscosity of the material, and so calculations with melt would allow for both shear and bulk attenuation phenomenon to be explored. However, perhaps most interesting would be to consider the temperature dependent properties of the grain boundaries and the triple lines and explore the effect of premelting and sub-solidus mechanisms of attenuation more generally.
6 Conclusions
We have presented a simple grain-scale model of diffusionally-accommodated/-assisted grain-boundary sliding and calculated its macroscale viscoelastic rheology. The rheology can be well-described by an extended Burgers model with a power-law distribution of relaxation times and a long period cut-off. The parameters of this extended Burgers model are in remarkable agreement with those inferred from laboratory experiments on the rock-analogue material borneol. The Andrade exponent $\alpha$ is around a third, the Andrade time $\tau_{A}$ is comparable to the Maxwell time $\tau_{M}$, and the relaxation strength $\Delta$ is around 0.6. The close agreement between the laboratory experiments and the model gives confidence that the model can be used to describe the viscoelastic behaviour of a wide range of polycrystalline materials in the regimes of diffusionally-accommodated/-assisted grain-boundary sliding. However, laboratory experiments also indicate that there remain additional mechanisms of attenuation that are still to be fully understood, particularly as the temperature of the material approaches the melting point.
Appendix A Governing equations
A.1 Elasticity
Grain interiors are assumed to behave linear elastically, with conservation of momentum
$\nabla\cdot\underline{\underline{\boldsymbol{\sigma}}}=\mathbf{0}$ (11)
for a symmetric stress tensor $\underline{\underline{\boldsymbol{\sigma}}}$, related to the displacement $\mathbf{u}$ by the constitutive law
$\sigma_{ij}=\lambda e_{kk}\delta_{ij}+2\mu e_{ij}$ (12)
where $\lambda$ and $\mu$ are the Lame moduli and $\underline{\underline{\boldsymbol{\mathrm{e}}}}$ is the symmetric strain tensor,
$e_{ij}=\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right).$ (13)
A.2 Grain boundary diffusion
On the boundaries between grains the elastic displacement is discontinuous. Diffusion of atoms and vacancies along the grain boundaries leads to plating out or removal of material. Grain boundary diffusion is described by [e.g. 7, section 2.2]
$\left[\dot{\mathbf{u}}\cdot{\mathbf{n}}\right]+\Omega\delta D^{\text{gb}}_{\text{v}}\nabla_{\perp}^{2}c=0$ (14)
where square brackets are used to denote the jump in a quantity across the grain boundary. $\dot{\mathbf{u}}\equiv\partial\mathbf{u}/\partial t$ is the velocity, the rate of change of the displacement. ${\mathbf{n}}$ is a unit vector normal to the grain boundary. $\left[\dot{\mathbf{u}}\cdot{\mathbf{n}}\right]$ represents the jump in normal velocity across the grain boundary, which is the rate at which new material is plated out on the grain boundary. This plating is accommodated by diffusion, where $D^{\text{gb}}_{\text{v}}$ is the vacancy diffusivity along the grain boundary, $\delta$ is the grain boundary width, $\Omega$ is the atomic volume, and $c$ is the concentration of vacancies. $\nabla_{\perp}$ represents the gradient operator restricted to the grain boundary; $\nabla_{\perp}^{2}$ is the surface Laplacian operator.
The concentration of vacancies is related to the normal stresses on the boundary by Herring’s relation [36]
$c=c_{0}\left(1+\frac{\Omega p}{kT}\right)$ (15)
where $c_{0}$ is the equilibrium concentration of vacancies, $k$ is Boltzmann’s constant and $p$ is the normal component of the traction on the boundary,
$p\equiv{\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\sigma}}}\cdot{\mathbf{n}}\quad\text{on }S,$ (16)
where $S$ is the grain-boundary interface. (15) can be substituted into (14) to write
$\left[\dot{\mathbf{u}}\cdot{\mathbf{n}}\right]+\frac{\Omega\delta D^{\text{gb}}}{kT}\nabla_{\perp}^{2}p=0,\$ (17)
where $D^{\text{gb}}=D^{\text{gb}}_{\text{v}}c_{0}$ is the self-diffusion coefficient for grain boundary diffusion.
A.3 Grain boundary sliding
The simplest model of grain boundary sliding is to consider the interface between two grains as if it were coated with a viscous fluid that offers resistance to being sheared,
${\mathbf{n}}\times\underline{\underline{\boldsymbol{\sigma}}}\cdot{\mathbf{n}}=\frac{\eta_{\text{gb}}}{\delta}[{\mathbf{n}}\times\dot{\mathbf{u}}]\quad\text{ on }S$ (18)
where $\eta_{\text{gb}}$ is the grain boundary viscosity and $\delta$ is the grain boundary width. In the present work, we focus on timescales long enough such that the shear stresses on the grain boundaries are relaxed, and thus impose
[…]
Together (11), (16), (17) and (19) describe a coupled problem to be solved for the displacement $\mathbf{u}$ inside the grains and the normal stresses $p$ on the interfaces between grains.
Appendix B Laplace transform
A standard approach to studying the viscoelastic response of a material is to Laplace transform the variables in time as
$\tilde{f}(s)=\int_{0}^{\infty}f(t)\mathrm{e}^{-st}\;\mathrm{d}t$ (20)
where tildes are used to denote Laplace-transformed variables. Under this transform differentiation in time becomes multiplication by $s$, assuming initial conditions are such that all variables vanish before time 0 i.e. before the application of some forcing (an assumption of causality). Laplace transforming replaces $\dot{\mathbf{u}}$ by $s\tilde{\mathbf{u}}$. In the sections that follow we work with Laplace transformed variables, but for clarity of notation we do not explicitly write tildes above each Laplace-transformed variable.
Appendix C Weak forms
For finite element analysis the governing equations are most usefully written in weak form. Consider a domain $V$ potentially containing many individual grains. Taking the dot product of the complex conjugate of (11) with a vector test function $\mathbf{v}$, integrating over the domain $V$ and using Green’s identity yields
$\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{i}}p^{*}[\mathbf{v}\cdot{\mathbf{n}}]\;\mathrm{d}S=\int_{S_{0}}{\mathbf{t}}^{0*}\cdot\mathbf{v}\;\mathrm{d}S$ (21)
where $S_{i}$ represents grain boundary surfaces; $S_{0}$ represents any exterior surfaces; and ${\mathbf{t}}^{0}$ any imposed tractions on those exterior surfaces. $\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}}$ is the stress tensor formed from the displacement $\mathbf{u}$. $\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}$ is the symmetric strain tensor formed from the test function $\mathbf{v}$. The colon (::::) denotes the tensor contraction on two indices. The terms involving the normal and tangential components of the traction at grain boundaries have been rewritten using (16) and (19). Superscript denotes complex conjugation. Note that both the displacement and the test function are complex functions, and the Laplace transform parameter $s$ may also be complex.
Multiplying the complex conjugate of (17) with a scalar test function $q$, integrating over the grain boundary surfaces and using Green’s identity yields
$-\int_{S_{i}}[\mathbf{u}^{*}\cdot{\mathbf{n}}]q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S=-\frac{\Omega\delta}{s^{*}}\int_{C}{\mathbf{j}_{\text{v}}^{*0}}\cdot\boldsymbol{\nu}\,q\;\mathrm{d}l$ (22)
where $C$ represents the triple lines where grain boundary surfaces intersect and $\boldsymbol{\nu}$ is an outward-pointing conormal at the triple line (i.e. a vector perpendicular to both the grain-boundary normal and a vector in the direction of the contact line). $\mathbf{j}^{0}_{\text{v}}$ is the flux of vacancies into the triple line. It will be assumed that the integral over $C$ vanishes, which is an assumption that there is no net flux into or out of the triple lines.
Appendix D Periodic homogenisation
Consider a collection of identical grains which tessellate space. The displacement within each grain can be decomposed as
$\mathbf{u}=\mathbf{U}+\hat{\mathbf{u}}$ (23)
where $\mathbf{U}$ represents a rigid body motion and $\hat{\mathbf{u}}$ represents the remaining elastic deformation. Each grain in the tessellation will be assumed to have the same elastic deformation $\hat{\mathbf{u}}$, but that $\mathbf{U}$ will vary in a systematic and prescribed manner. Following [8] (section 3), the kinematics of the rigid body motions will be prescribed in terms of macroscale strain measures $\underline{\underline{\boldsymbol{\Gamma}}}$ and $\underline{\underline{\boldsymbol{K}}}$, where the symmetric part of $\underline{\underline{\boldsymbol{\Gamma}}}$ represents the macroscale strain tensor. The jump in the rigid body displacement across the grain boundaries is [8, equation (21)]
$\left[\mathbf{U}\right]=\underline{\underline{\boldsymbol{\Gamma}}}\cdot\mathbf{R}+\left(\underline{\underline{\boldsymbol{K}}}\cdot\mathbf{R}\right)\times{\mathbf{d}}$ (24)
where $\mathbf{R}$ is a vector which joins the centroids of the neighbouring grains, and ${\mathbf{d}}$ is a vector joining the centroid of the neighbouring grain to the position on the grain boundary [8, Figure 1]. In the present work we will only consider upscaling to a Cauchy continuum, not a more general micropolar continuum as was done in [8]. We will thus neglect the strain measure $\underline{\underline{\boldsymbol{K}}}$ associated with gradients in microrotation, prescribing
$\left[\mathbf{U}\cdot{\mathbf{n}}\right]={\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}\cdot\mathbf{R},$ (25)
and solve for a vector displacement field $\hat{\mathbf{u}}$ which is identical in each grain in the periodic tessellation. Substitution of (23) and (25) into the weak forms (21) and (22) yields
$\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\hat{\mathbf{u}}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{i}}p^{*}[\mathbf{v}\cdot{\mathbf{n}}]\;\mathrm{d}S=0$ (26)
$-\int_{S_{i}}[\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}]q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S=\int_{S_{i}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (27)
which given $\underline{\underline{\boldsymbol{\Gamma}}}$ can be solved for $\hat{\mathbf{u}}$ and $p$.
D.1 Equivalent energetics
In upscaling, we wish to replace the discrete collection of grains by an equivalent continuum description, with a macroscale effective stress tensor $\overline{\underline{\underline{\boldsymbol{\sigma}}}}$ related to the macroscale strain measure $\underline{\underline{\boldsymbol{\Gamma}}}$ by
$\overline{\sigma}_{ij}=C_{ijkl}\Gamma_{kl}$ (28)
We determine the effective stiffness tensor $C_{ijkl}$ by demanding that both the grain-scale model and the upscaled continuum have the same energetics, i.e. they store and dissipate energy at the same rate [8]. This is achieved by demanding equality between the microscale
$\Psi=\frac{1}{V}\left(\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{u}}\;\mathrm{d}V-\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}p\;\mathrm{d}S\right)$ (29)
and the macroscale
$\Psi=\overline{\sigma}_{ij}\Gamma_{ij}=\Gamma_{ij}C_{ijkl}^{*}\Gamma_{kl}^{*}.$ (30)
Appendix E Scaling
The equations can be simplified further by scaling. There is a natural lengthscale associated with the size of the grains, which we label $d$. If $u_{0}$ is a typical displacement, then a natural scale for stresses is $\mu u_{0}/d$ where $\mu$ is the intrinsic elastic shear modulus of the grains. Importantly, there is a natural timescale, given by
$\tau=\frac{kTd^{3}}{\delta D^{\text{gb}}\Omega\mu}$ (31)
which up to a numerical factor is a Maxwell time for the material, a ratio of the steady-state Coble creep viscosity to the shear modulus. Once the governing equations have been scaled, there are only two dimensionless parameters which remain to describe the rheology: the dimensionless Poisson’s ratio $\nu$ and the dimensionless Laplace transform parameter $s\tau$ (or equivalently the dimensionless frequency $\omega\tau$ of oscillation when considering time harmonic loading).
Appendix F Numerical Implementation
The computational domain consists of a single grain. A finite element mesh of the domain is constructed (triangular in 2D, tetrahedral in 3D). The variable $\hat{\mathbf{u}}$ is solved for throughout the domain, the variable $p$ solved for only on the exterior boundary of the domain. Periodic boundary conditions on $p$ are applied to impose invariance under translations by lattice vectors, where a point on each grain boundary is identical under translation to a point on the opposite face. Since the grain boundaries are on the edges of the computational domain, each section of domain boundary contains half the total grain boundary, but by periodicity the same boundary is traversed twice when integrating round the total surface of the grain. This simplifies the calculation of quantities involving jumps across the interface, as a single integral on a grain boundary involving a jump can be split into two separate integrals only involving displacements on one side of the boundary. The weak forms (32) and (33) can then be written as
$\int_{V_{\text{grain}}}\underline{\underline{\boldsymbol{\sigma}}}^{\hat{\mathbf{u}}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{\text{grain}}}p^{*}\mathbf{v}\cdot{\mathbf{n}}\;\mathrm{d}S$ $=0$ (32)
$-\int_{S_{\text{grain}}}\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}\,q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{2kTs^{*}}\int_{S_{\text{grain}}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S$ $=\frac{1}{2}\int_{S_{\text{grain}}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (33)
where the factors of $1/2$ come from accounting for traversing the same boundaries twice as one integrates round a whole grain.
The weak form was discretised using a P2 vector element for $\hat{\mathbf{u}}$ and a P2 scalar element for $p$. The resulting set of linear equations were solved iteratively using MINRES and a GAMG preconditioner. The weak forms were discretised using the dolfinx python library [37] with linear algebra performed using PETSc [38].
Figure 8: Influence of mesh resolution on finite element calculation of attenuation. The attenuation $Q^{-1}$ is plotted against the normalized frequency $\omega\tau_{M}$ for hexagonal grains for different values of $n$, the number of mesh cells on each edge of the hexagon.
The accuracy and precision of the numerical results depends on mesh resolution. The effect of changing resolution is illustrated in Figure 8 which shows calculations of attenuation for hexagonal grains. The mesh of the regular hexagon is made up of cells of equilateral triangles of equal size, and the resolution specified by a parameter $n$ which gives the number of mesh cells per edge. Increasing $n$ by a factor of 2 increases the total number of mesh cells by a factor of 4. Simulations shown in the main text use $n=256$ which has 393,216 triangular mesh cells. As can be seen in Figure 8 mesh resolution limits the highest frequencies that can be resolved, where the attenuation is underestimated at high frequencies when the resolution is too coarse. This is as expected from equation (5), where diffusion acts over a length scale $l\propto\omega^{-1/3}$ at high frequencies, and this scale must be well-resolved to get accurate estimates of attenuation. Each doubling of $n$ leads to an eightfold-increase in frequency that can be resolved.
Producing well-resolved simulations in three-dimensions is computationally much more demanding than in two-dimensions. The simulations here were performed at a range of resolutions up to $n=32$ mesh cells per edge of the tetrakaidecadedron, with a total of $2,033,250$ tetrahedral mesh cells. Results shown in the main text (Figure 4) are limited to frequencies lower that $\omega\tau_{M}<2\times 10^{3}$ to be restricted to a range of frequencies that are well resolved.
Appendix G Viscoelastic response functions
The main outcome of the finite element modelling is the tensor $C_{ijkl}$ which relates macroscale stress and strain in the Laplace-transformed variables of (28). For the simplest cases, this response is isotropic and can be written
$C_{ijkl}=s\tilde{K}(s)\delta_{ij}\delta_{kl}+s\tilde{G}(s)\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\frac{2}{N}\delta_{ij}\delta_{kl}\right)$ (34)
where $N$ is the dimension of the space, and $\tilde{K}(s)$ and $\tilde{G}(s)$ are the Laplace transforms of the bulk and shear relaxation moduli. The factors of $s$ in the definitions above arise from the standard definition of $G(t)$ in the time domain as the stress response to a unit step of strain. Knowledge of the functions $\tilde{K}(s)$ and $\tilde{G}(s)$ completely describes the response of the material, but depending on the type of loading considered other forms of response function may be preferred. For example, the strain response to a unit step of stress is termed the creep compliance and denoted $J(t)$ in the time domain. The Laplace-transformed variables satisfy the reciprocal relation
$s\tilde{J}(s)=\frac{1}{s\tilde{G}(s)}.$ (35)
Detailed discussion of the conversions between different sets of viscoelastic response function can be found in textbooks (e.g. [39, 40]).
One of the most common experimental tests of viscoelastic behaviour is a forced oscillation test, where the sample has a time-harmonic deformation of a given frequency $\omega$ imposed. The stress response to a unit time harmonic strain is the complex shear modulus, given by
$G^{*}(\omega)=\left.s\tilde{G}(s)\right|_{s=i\omega}.$ (36)
The corresponding strain response to unit time harmonic stress is the complex shear compliance, given by
$J^{*}(\omega)=\left.s\tilde{J}(s)\right|_{s=i\omega}$ (37)
where there is the reciprocity $J^{*}(\omega)G^{*}(\omega)=1$. The superscript * on $J^{*}(\omega)$ and $G^{*}(\omega)$ does not denote complex conjugation, but is a standard notation used for these quantities. As complex quantities, $G^{*}(\omega)$ and $J^{*}(\omega)$ can be written in terms of real and imaginary parts as
$G^{*}(\omega)$ $=G_{1}(\omega)+iG_{2}(\omega),$ (38)
$J^{*}(\omega)$ $=J_{1}(\omega)-iJ_{2}(\omega).$ (39)
The reciprocal quality factor or loss tangent is
$Q^{-1}(\omega)=\tan\arg G^{*}(\omega)=\frac{G_{2}(\omega)}{G_{1}(\omega)}=\frac{J_{2}(\omega)}{J_{1}(\omega)}.$ (40)
Appendix H Response functions for simple models of viscoelasticity
For reference we provide here the response functions for some standard models of viscoelasticity. In each case $J_{0}$ refers to the unrelaxed compliance.
H.1 Maxwell
The Maxwell model is represented by a spring and dashpot connected in series and has compliance
$s\tilde{J}(s)=J_{0}\left(1+\frac{1}{s\tau_{M}}\right)$ (41)
where $\tau_{M}$ is the Maxwell time. In the time domain the corresponding creep function is
$J(t)=J_{0}\left(1+\frac{t}{\tau_{M}}\right).$ (42)
H.2 Kelvin-Voigt
The Kelvin-Voigt model is represented by a spring and dashpot connected in parallel and has compliance and creep function
$s\tilde{J}(s)$ $=\frac{J_{r}}{1+s\tau_{r}},$ (43)
$J(t)$ $=J_{r}\left(1-\mathrm{e}^{-t/\tau_{r}}\right),$ (44)
where $\tau_{r}$ is the relaxation time and $J_{r}$ is the relaxed compliance.
H.3 Burgers
The Burgers model combines a Maxwell model and a Kelvin-Voigt model in series, with compliance and creep function
$s\tilde{J}(s)$ $=J_{0}\left({1+\frac{\Delta}{1+s\tau_{r}}+\frac{1}{s\tau_{M}}}\right),$ (45)
$J(t)$ $=J_{0}\left(1+\Delta\left(1-\mathrm{e}^{-t/\tau_{r}}\right)+\frac{t}{\tau_{M}}\right),$ (46)
where $\Delta$ is the anelastic relaxation strength.
H.4 Andrade
The Andrade model has compliance and creep function
$s\tilde{J}(s)$ $=J_{0}\left({1+\frac{\Gamma(1+\alpha)}{(s\tau_{A})^{\alpha}}+\frac{1}{s\tau_{M}}}\right),$ (47)
$J(t)$ $=J_{0}\left(1+\left(\frac{t}{\tau_{A}}\right)^{\alpha}+\frac{t}{\tau_{M}}\right),$ (48)
where $\alpha$ is a dimensionless Andrade exponent, $\tau_{A}$ is the Andrade timescale and $\Gamma(z)$ is the Gamma function.
H.5 Extended Burgers
The extended Burgers model has
$s\tilde{J}(s)$ $=J_{0}\left({1+\Delta\int_{0}^{\infty}\frac{\gamma\rho(\gamma)}{s+\gamma}\;\mathrm{d}\gamma+\frac{1}{s\tau_{M}}}\right),$ (49)
$J(t)$ $=J_{0}\left({1+\Delta\int_{0}^{\infty}\left(1-\mathrm{e}^{-\gamma t}\right)\rho(\gamma)\;\mathrm{d}\gamma+\frac{t}{\tau_{M}}}\right),$ (50)
for a density function (relaxation spectrum) $\rho(\gamma)$ as a function of frequency, which satisfies $\rho(\gamma)\geq 0$ and
$\int_{0}^{\infty}\rho(\gamma)\;\mathrm{d}\gamma=1.$ (51)
The Burgers model in (45) is a special case of (49) with a single frequency, $\rho(\gamma)=\delta(\gamma-1/\tau_{r})$, where $\delta(x)$ is the Dirac delta function. The integral in (49) can be recognised as the Stieltjes transform of $\gamma\rho(\gamma)$.
The extended Burgers model in (49) is sufficiently general to describe any linear viscoelastic response. A commonly-used parametrisation of $\rho(\gamma)$ has a power-law distribution in a period band between a short-period cut-off $\tau_{L}$ and a long-period cut-off $\tau_{H}$ (e.g. [41, 19]). In the present work we will neglect the short-period cut-off, and write the density as
$\rho(\gamma)=\begin{cases}\dfrac{\alpha\tau_{H}}{(\gamma\tau_{H})^{\alpha+1}},\quad\quad&\gamma\tau_{H}>1,\\ 0,\quad\quad&\text{otherwise.}\end{cases}$ (52)
in terms of the power law exponent $\alpha$. The Stieltjes transform can be determined analytically, and (49) can be written as
$s\tilde{J}(s)=J_{0}\left({1+\Delta\,{}_{2}F_{1}\left(1,\alpha,1+\alpha,-s\tau_{H}\right)+\frac{1}{s\tau_{M}}}\right)$ (53)
where ${}_{2}F_{1}$ is the ordinary hypergeometric function. The corresponding creep function (plotted in Figure 9) is
$J(t)=J_{0}\left(1+\Delta\left(1-\alpha\,E_{1+\alpha}\left(\frac{t}{\tau_{H}}\right)\right)+\frac{t}{\tau_{M}}\right)$ (54)
where $E_{n}(x)$ is the generalized exponential integral.
Figure 9: Scaled creep function $J(t)/J_{0}$ as a function of dimensionless time $t/\tau_{M}$ for an extended Burgers model with the parameters given in Table 1 for tetrakaidecahedral grains ($\alpha=\tfrac{1}{3}$, $\tau_{A}/\tau_{M}=1.68$, $\Delta=0.552$). The short-time behaviour is well-described by an Andrade model and the long-time behaviour by a Maxwell model.
Note that as $s\tau_{H}\rightarrow\infty$ (high frequencies)
${}_{2}F_{1}\left(1,\alpha,1+\alpha,-s\tau_{H}\right)\sim\frac{\Gamma(1+\alpha)\Gamma(1-\alpha)}{(s\tau_{H})^{\alpha}}$ (55)
and so the extended Burgers model behaves in the same way as the Andrade model at high frequencies, with parameters related by
$\tau_{A}=\tau_{H}\left(\Gamma(1-\alpha)\,\Delta\right)^{-1/\alpha}.$ (56)
As $s\tau_{H}\rightarrow 0$ (low frequencies),
$s\tilde{J}(s)\sim J_{0}\left({1+\Delta+\frac{1}{s\tau_{M}}}\right)$ (57)
and the extended Burgers model behaves as a Maxwell material.
Appendix I Voigt averaging
Figures 4 and 5 show results for tetrakaidecahedral grains with Voigt-averaged properties. The Voigt-averaged shear modulus is given in the Laplace domain by
$s\tilde{G}^{V}(s)=\frac{2}{5}s\tilde{G}^{(1)}(s)+\frac{3}{5}s\tilde{G}^{(2)}(s),$ (58)
where superscript $V$ denotes the Voigt-average, and superscripts (1) and (2) represent deformation with different principal axes: (1) has principal axes aligned with the square faces (see equation (40) of [7]). It should be noted that the quantities presented in Figure 4 average in slightly different ways, although the averaging formulae are a direct consequence of the above expression. For example, for $s\ll 1$ the material behaves as a Maxwell model for which
$s\tilde{G}(s)\sim s\eta-\frac{s^{2}\eta^{2}}{G_{ss}}+\cdots$ (59)
and hence the steady-state viscosity is Voigt-averaged as
$\eta^{V}=\frac{2}{5}\eta^{(1)}+\frac{3}{5}\eta^{(2)}$ (60)
whereas the steady-state modulus is Voigt-averaged as
$\frac{(\eta^{V})^{2}}{G_{ss}^{V}}=\frac{2}{5}\frac{(\eta^{(1)})^{2}}{G_{ss}^{(1)}}+\frac{3}{5}\frac{(\eta^{(2)})^{2}}{G_{ss}^{(2)}}.$ (61)
Similarly, for $s\gg 1$ the material behaves as an Andrade model with
$s\tilde{G}(s)\sim G_{0}-\frac{G_{0}\Gamma(1+\alpha)}{\left(s\tau_{A}\right)^{\alpha}}+\cdots$ (62)
and thus the unrelaxed modulus averages as
$G_{0}^{V}=\frac{2}{5}G_{0}^{(1)}+\frac{3}{5}G_{0}^{(2)},$ (63)
while the Andrade time averages as
$\frac{G_{0}^{V}}{\left(\tau_{A}^{V}\right)^{\alpha}}=\frac{2}{5}\frac{G_{0}^{(1)}}{\left(\tau_{A}^{(1)}\right)^{\alpha}}+\frac{3}{5}\frac{G_{0}^{(2)}}{\left(\tau_{A}^{(2)}\right)^{\alpha}}.$ (64)
It should also be noted that the Voigt-averaged Maxwell time is defined by $\tau_{M}^{V}=\eta^{V}/G_{0}^{V}$.
Appendix J Rational fits to functions of Poisson’s ratio
The data shown in Figures 1 and 4 as a function of Poisson’s ratio $\nu$ can be effectively parametrised by fitting a rational function of the form
$r(\nu)=\frac{\displaystyle\sum_{j=1}^{m}\frac{w_{j}f_{j}}{\nu-\nu_{j}}}{\displaystyle\sum_{j=1}^{m}\frac{w_{j}}{\nu-\nu_{j}}}$ (65)
where $\nu_{j}$ are the nodes, $f_{j}$ are function values, and $w_{j}$ are weights. The rational function has the property that $r(\nu_{j})=f_{j}$. The above equation expresses the rational function in barycentric form, and an alternative and equivalent expression is
$r(\nu)=\frac{\displaystyle\sum_{j=1}^{m}w_{j}f_{j}\prod_{k\neq j}(\nu-\nu_{k})}{\displaystyle\sum_{j=1}^{m}w_{j}\prod_{k\neq j}(\nu-\nu_{k})},$ (66)
from which it is clear the function is the ratio of two polynomials of degree $m-1$. The weights can all be scaled by a constant factor and the rational function will be unchanged. To specify the weights uniquely, here we fix $w_{1}=1$.
$f_{1}$ $f_{2}$ $w_{2}$
$G_{0}/\mu$ 0.599973 0.857133 -0.700034
$G_{ss}/\mu$ 0.353374 0.686123 -0.515031
$(\tau_{A}/\tau_{M})^{\alpha}$ 0.68953 1.319737 -0.301846
Table 2: Parameters of a rational fit to the data shown in Figure 1 as a function of $\nu$ for hexagonal grains. In each case $\nu_{1}=-1$, $\nu_{2}=\tfrac{1}{2}$, and $w_{1}=1$. The Andrade exponent $\alpha=0.3672092$.
Table 2 gives parameters which fit the data for hexagons shown in Figure 1. The fits have just three free parameters, and represent the properties as a ratio of two linear functions of $\nu$ (i.e. $m=2$). The fits for the scaled moduli $G_{0}/\mu$ and $G_{ss}/\mu$ have relative errors $<10^{-6}$ which suggests that the representation of these quantities in terms of a ratio of linear functions may in fact be exact, similar to the analytic result for $G_{0}/\mu$ for spheres by Zener [42]. Larger errors are seen in fitting the scaled Andrade time $\tau_{A}/\tau_{M}$, but the parameters in the table still represent an excellent fit with relative errors $<10^{-3}$.
$f_{1}$ $f_{2}$ $f_{3}$ $w_{2}$ $w_{3}$
$G^{(1)}_{0}/\mu$ 0.249839 0.74434 0.556733 1.404701 -2.491113
$G^{(2)}_{0}/\mu$ 0.611883 0.873231 -0.791083
$G^{V}_{0}/\mu$ 0.46641 0.821582 0.685598 0.97236 -1.981183
$G^{(1)}_{ss}/\mu$ 0.029394 0.498677 0.265034 1.428164 -2.599101
$G^{(2)}_{ss}/\mu$ 0.135534 0.693088 0.491729 2.490243 -3.739137
$G^{V}_{ss}/\mu$ 0.089789 0.56882 0.390075 2.503946 -3.76036
$(\tau^{(1)}_{A}/\tau_{M}^{(1)})^{\alpha}$ 0.188561 1.176442 0.696462 22.349757 -32.371228
$(\tau^{(2)}_{A}/\tau_{M}^{(2)})^{\alpha}$ 0.35751 1.862253 1.123438 1.583196 -2.715582
$(\tau^{V}_{A}/\tau_{M}^{V})^{\alpha}$ 0.308609 1.444995 0.901722 10.238062 -14.343569
Table 3: Parameters of a rational fit to the data shown in Figure 4 as a function of $\nu$ for tetrakaidecahedral grains. In each case $\nu_{1}=-1$, $\nu_{2}=\tfrac{1}{2}$, $\nu_{3}=0$, and $w_{1}=1$. The Andrade exponent $\alpha=\tfrac{1}{3}$.
Table 3 gives parameters for fits of the data plotted in Figure 4 for tetrakaidecahedral grains. Unlike the case of hexagons, the data require higher order polynomials to be fit well; the table uses a fit of a ratio of quadratics ($m=3$), which fits the scaled moduli to within a relative error of $<10^{-3}$, and the scaled Andrade times $\tau_{A}/\tau_{M}$ to within $<2\times 10^{-2}$.
Appendix K Scaling argument for high frequency attenuation
The scaling argument given here follows that in Lee et al. [6]. For material deforming at a typical strain $\epsilon$, the energy stored elastically per unit volume scales as
$\mathcal{U}\sim\mu\epsilon^{2}.$ (67)
The energy dissipated by grain boundary diffusion over a cycle scales as
$\mathcal{D}\sim\frac{\Omega\delta D^{\text{gb}}}{kT\omega V}\int_{S_{i}}|\nabla_{\perp}p|^{2}\;\mathrm{d}S=\frac{l^{3}}{\mu V}\int_{S_{i}}|\nabla_{\perp}p|^{2}\;\mathrm{d}S$ (68)
where $l$ is the lengthscale defined in (5).
With a stress singularity at the triple junction, the normal stress scales as
$p\sim\mu\epsilon\left(r/d\right)^{-\lambda}$ (69)
where $d$ is a measure of grain size, and $r$ is distance from the triple junction. Thus the gradient scales as
$\nabla_{\perp}p\sim\frac{\mu\epsilon}{d}\left(r/d\right)^{-(\lambda+1)}.$ (70)
The surface integral term in (68) can be approximated in 2D as
$\int_{l}^{\infty}|\nabla_{\perp}p|^{2}\,\mathrm{d}r\sim\frac{\mu^{2}\epsilon^{2}}{d}\left(l/d\right)^{-(2\lambda+1)}$ (71)
where it is assumed that diffusion smooths the singularity within a distance $l$ of the triple junction, and that $l\ll d$. In 2D the grain volume $V\sim d^{2}$ and hence (68) becomes
$\mathcal{D}\sim\mu\epsilon^{2}\left(l/d\right)^{2(1-\lambda)}\sim\mu\epsilon^{2}\left(\omega\tau_{M}\right)^{-2(1-\lambda)/3}.$ (72)
The attenuation can be related to the ratio of the energy dissipated over a cycle to the energy stored and hence
$Q^{-1}\sim\frac{\mathcal{D}}{\mathcal{U}}\sim\left(\omega\tau_{M}\right)^{-\alpha}$ (73)
where
$\alpha=\frac{2}{3}\left(1-\lambda\right).$ (74)
Appendix L Volume diffusion
The main focus of this work has been on the process of grain boundary diffusion, as that process has been suggested to be the one of relevance in a number of experimental studies [e.g. 2]. However, there are circumstances where diffusion within the interiors of grains dominates over diffusion on the grain boundaries, and in steady-state this is known as Nabarro-Herring creep. From a modelling perspective volume diffusion can be treated in a very similar way to surface diffusion, and simply results in replacing a surface integral in the weak form of the equations by a volume integral, i.e. instead of (33), one has
$-\int_{S_{\text{grain}}}\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}\,q\;\mathrm{d}S+\frac{\Omega D}{kTs^{*}}\int_{V_{\text{grain}}}\nabla p^{*}\cdot\nabla q\;\mathrm{d}V=\frac{1}{2}\int_{S_{\text{grain}}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (75)
where $D$ is the self-diffusion coefficient for volume diffusion (see [7, 8]).
Figure 10: Plots similar to Figure 2 for hexagonal grains with volume diffusion rather than surface diffusion. The extended Burgers fit has $\Delta=0.281$, $\alpha=0.5508138$ and $\tau_{A}/\tau_{M}=0.99$.
Figure 10 illustrates the viscoelastic behaviour associated with volume diffusion for the case of hexagons. The results are again scaled using a Maxwell time $\tau_{M}=\eta/G_{0}$ where $G_{0}$ is exactly the same as before (as the pure elastic problem is unchanged) but the steady-state viscosity is different and given by
$\eta=\frac{5}{288}\frac{kTd^{2}}{D\Omega},$ (76)
proportional to the grain size squared rather than cubed. The above expression has a different numerical prefactor to the expression for hexagons given in equation (36) of [7] (a prefactor of 1/36). This prefactor difference is a consequence of the difference in homogenisation scheme, with the scheme used here being the same as the equivalent dissipation principle in [8].
The principal difference in viscoelastic behaviour between grain boundary diffusion and volume diffusion concerns the Andrade exponent $\alpha$, which can be understood using a similar scaling argument to that presented in Appendix K. The length scale over which volume diffusion acts when forced at a frequency $\omega$ is
$l=\left(\frac{\mu D\Omega}{kT\omega}\right)^{1/2},$ (77)
which replaces the surface diffusion result in (5). The expression for dissipation in (68) becomes
$\mathcal{D}\sim\frac{\Omega D}{kT\omega V}\int_{V}|\nabla p|^{2}\;\mathrm{d}V=\frac{l^{2}}{\mu V}\int_{V}|\nabla p|^{2}\;\mathrm{d}V$ (78)
and the expression in (72) becomes
$\mathcal{D}\sim\mu\epsilon^{2}\left(l/d\right)^{2(1-\lambda)}\sim\mu\epsilon^{2}\left(\omega\tau_{M}\right)^{-(1-\lambda)},$ (79)
and hence the expression relating the Andrade exponent to the stress exponent in (74) becomes
$\alpha=1-\lambda.$ (80)
Thus on attenuation plots such as Figure 10b volume diffusion shows a steeper slope than grain boundary diffusion does at high frequencies (steeper by a factor of 3/2).
Acknowledgments
I thank H. Yamauchi and Y. Takei for kindly sharing their data. I also thank H. Innes, G. Johnson, T. Breithaupt, D. Al-Attar, Y. Takei and D. McKenzie for insightful discussions on transient creep.
Data accessibility
Python code used for the finite element simulations is available at https://github.com/johnrudge/transient_diffusion_creep. The results of the finite element calculations (used in plotting the figures) can be found in a spreadsheet in the supplemental information.
Funding statement
This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.